A firing pin pushes a bullet of 0.8kg with a force of 2000 N from rest out the chamber in less than 0.1 seconds. What is the velocity in which it leaves the chamber ?
1
Expert's answer
2018-05-27T08:40:08-0400
Let’s first find the acceleration of the bullet from the Newton’s Second Law of Motion:
F=ma,
here, F is the force with which the firing pin pushes the bullet, m is the mass of the bullet and a is the acceleration of the bullet.
Then, from this formula we can find the acceleration of the bullet:
a=F/m=(2000 N)/(0.8 kg)=2500 m/s^2 .
We can find the velocity of the bullet as it leaves the chamber from the kinematic equation:
v=v_0+at,
here, v_0=0 is the initial velocity of the bullet, v is the velocity of the bullet as it leaves the chamber (or the final velocity), a is the acceleration of the bullet and t is the time that the bullet needs to leave the chamber.
Then, from this formula we can find the final velocity of the bullet:
v=at=2500 m/s^2 ∙0.1 s=250 m/s.